# Active Gel Theory for Cell Migration With Two Myosin Isoforms

### v1

@@ III.5 The Rigid Limit and the State Diagram
We now neglect any dynamics of the background stress by assuming $\delta s_{0}(t)\equiv 0$. From the stress boundary condition, we can infer $s^{0}=0$. The equilibrium values for the isoforms are chosen, as before, to be $c_{\text{A}}^{0}=1$ and $c_{\text{B}}^{0}=c_{\text{A}}^{0}/c_{r}$. When solving Eqs. (12) numerically using the FVM, we however keep the dependence on the background stress. It then has to be determined for every time step, such that all boundary conditions are satisfied. In the analytical treatment, the perturbations are decomposed into Fourier modes of the form $\left[X_{i}(q_{i})\cos(q_{i}u)+Y_{i}(q_{i})\sin(q_{i}u)\right]e^{\sigma_{q}t}$ that have to satisfy the boundary conditions, leading to
@@
Figure 5: Motility modes in the rigid limit. (a) State diagram of the four different motility modes depending on the relative diffusion $K_{r}$ vs. the rescaled Péclet number $\lambda$. Numerical results obtained by FVM simulations of equations (12) are coded in levels of gray. The curves are obtained from the linear stability eigenvalue problem and mark analytically predicted borders of the state diagram. (b) Kymographs for the two motor isoform concentrations and the resulting total active stress in a state of stable oscillations ($\lambda=44$). (c) Examples of the four different motility modes as plots of velocity over time. $\lambda$ is varied while $K_{r}\approx 4$ was kept fixed. Parameters correspond to the points indicated in (a). Parameters not mentioned have the same values as in Fig. 3.
The results are shown in Fig. 5. As shown in panel (a), the numerical results of the full system of equations (12) – shown as background colors – agree very well with the linear stability analysis (curves). This a posteriori justifies the neglect of perturbations in $s_{0}$. We can identify four different modes of motility: sessile steady states, oscillations decaying towards a sessile steady state (transient oscillations), stable oscillations and motile steady states. Exemplary plots of velocity as a function of time for each of these modes are shown in panel (c).
@@
Note that in the linear analysis we have only used the first mode ($n=1$), as it is the most dominant one for the dynamics: in fact, it is the first antisymmetric mode that creates the strongest polarity for $\delta c_{A}$ and $\delta c_{B}$. Modes of even $n$ do not produce cell motility as they create symmetric concentration profiles. Higher modes of odd $n$ cause antisymmetric concentration profiles that consist of multiple maxima/minima, leading to overall less polarity. Hence, the first mode suffices for a qualitative analysis of the dynamics of the systems, as also confirmed by the numerical solution. Also note that performing an analogous linear stability analysis for the original system Eq. (10) by simply fixing the length does not yield results consistent with the corresponding FVM simulations.
@@ IV Discussion
The two-species model introduced here recovers the experimental distributions of the isoforms and in addition predicts the existence of a variety of migratory modes. Regarding the steady states, in the non-motile steady state both isoforms are homogeneously distributed and cause a symmetric active contractile stress. At sufficiently large Péclet numbers, however, this symmetry is broken and polarized concentration fields and, consequently, polarized active stress profiles emerge. This onset of polarization and migration is due to a supercritical pitchfork bifurcation. For a steadily moving cell beyond the bifurcation, we find the experimentally observed, isoform-specific myosin distribution profiles [12]: the faster NMIIA accumulates at the leading edge, while the stronger NMIIB localizes at the trailing edge.
@@
We have explored further modes of motility besides the steady state solutions. Oscillations occur in case that the more contractile isoform is also more diffusive (i.e. $P_{B}>P_{A},\ K_{r}>1$), via a cyclic pull-push-mechanism. This mechanism works best for similar total abundances of both isoform, $c_{r}\sim 1$, which might be the reason why it has not been observed yet. It has been suggested that different myosin isoforms can be regulated independently [52, 53, 54], which would allow for the relative diffusion to be controlled via $K_{r}$. Albeit not observed in the cell type investigated in Refs. [16, 12], it is not unlikely that these oscillations might occur in other cell types. They also could be induced by perturbations, either of the relative abundances or of the regulatory networks controlling actomyosin dynamics. In addition, the parameter $c_{r}$ could be tuned using optogenetical activation of knock-downs.
@@
In the oscillatory state, the amplitude of the cell length oscillation is much smaller than the one of velocity. This suggested to study the rigid limit, which enabled an analytic linear stability analysis and allowed to determine the full state diagram of the system, containing sessile cells, cells with transient oscillations and persistent oscillations, as well as steady motile cell solutions. This phase diagram is in good agreement with FVM simulations of the full model.
@@
In summary, here we have shown that active gel theory can be extended to also describe the effect of having two complementary isoforms at work. The fact that we could derive the same model from both kinetic and a thermodynamic perspectives demonstrate its fundamental nature and paves the way to also include other biologically relevant effects into the model. In particular, a third major molecular constituent of contractile structures in cells is $\alpha$-actinin, which like NMII also exists in different isoforms and which also competes for binding to the actin filaments. In the future, it would be interesting to further extend active gel theory to such situations with multiple components.

### v2

@@ III.5 The Rigid Limit and the State Diagram
We now neglect any dynamics of the background stress by assuming $\delta s_{0}(t)\equiv 0$. From the stress boundary condition, we can infer $s^{0}=0$. The equilibrium values for the isoforms are chosen, as before, to be $c_{\text{A}}^{0}=1$ and $c_{\text{B}}^{0}=c_{\text{A}}^{0}/c_{r}$. When solving Eqs. (12) numerically using the finite volume method, we however keep the dependence on the background stress. It then has to be determined for every time step, such that all boundary conditions are satisfied. In the analytical treatment, the perturbations are decomposed into Fourier modes of the form $\left[X_{i}(q_{i})\cos(q_{i}u)+Y_{i}(q_{i})\sin(q_{i}u)\right]e^{\sigma_{q}t}$ that have to satisfy the boundary conditions, leading to
@@
Figure 6: Motility modes in the rigid limit. (a) State diagram of the four different motility modes depending on the relative diffusion $K_{r}$ vs. the rescaled Péclet number $\lambda$. Numerical results obtained by finite volume simulations of equations (12) are coded in levels of gray. The curves are obtained from the linear stability eigenvalue problem and mark analytically predicted borders of the state diagram. (b) Kymographs for the two motor isoform concentrations and the resulting total active stress in a state of stable oscillations ($\lambda=44$). (c) Examples of the four different motility modes as plots of velocity over time. $\lambda$ is varied while $K_{r}\approx 4$ was kept fixed. Parameters correspond to the points indicated in (a). Parameters not mentioned have the same values as in Fig. 4.
The results are shown in Fig. 6. As shown in panel (a), the numerical results of the full system of equations (12) — shown as background colors —  agree very well with the linear stability analysis (curves). This a posteriori justifies the neglection of perturbations in $s_{0}$. We can identify four different modes of motility: sessile steady states, oscillations decaying towards a sessile steady state (transient oscillations), stable oscillations and motile steady states. Exemplary plots of velocity as a function of time for each of these modes are shown in panel (c).
@@
Note that in the linear analysis we have only used the first mode ($n=1$), as it is the most dominant one for the dynamics at the given hydrodynamic length of $\mathcal{L}\sim 1.25$: in fact, it is the first antisymmetric mode that creates the strongest polarity for $\delta c_{A}$ and $\delta c_{B}$. Modes of even $n$ do not produce cell motility as they create symmetric concentration profiles. Higher modes of odd $n$ cause antisymmetric concentration profiles that consist of multiple maxima/minima, leading to overall less polarity. Hence, the first mode suffices for a qualitative analysis of the dynamics of the systems, as also confirmed by the numerical solution. Also note that performing an analogous linear stability analysis for the original system Eq. (10) by simply fixing the length does not yield results consistent with the corresponding finite volume simulations.
III.6 Effect of Actin Polymerization
Figure 7: Bifurcation diagrams for different polymerization velocities $v_{p}$. As soon as $v_{p}\neq 0$ the symmetry of the supercritical pitchfork bifurcation is broken and imperfect bifurcations with three branches emerge. (a) Bifurcation diagram for positive $v_{p}$. (b) Bifurcation diagrams for negative $v_{p}$. Encircled points mark changes in stability. (c) Concentration profiles taken at $\mathcal{K}=120$ for polymerization velocity $v_{p}=0.01$ at the stable solution branch, the unstable branch and the biased motile solutions (from left to right). Parameters used: $\mathcal{L}^{2}=1.25,K_{r}=0.2,P_{A}=0.05,P_{B}~{}=~{}0.18,c_{\infty}~{}=~{}4,c_{r}=1.5$.
Contraction-driven motility emerges by an internal symmetry break as explained above. Uneven actin polymerization at the cell boundaries breaks front-rear symmetry by construction and allows cells to quickly respond to external cues. For one-species models, this effect of actin polymerization has been demonstrated before [22, 20, 21]. Although here our focus is on contractility, for completeness we have also performed a similar analysis for the two-species model presented here.
Polymerization can be incorporated into our model by modifying the force balance equation of Eq. 2 adding the constant polymerization velocity $v_{p}$ to the contraction driven velocity $v(l_{\pm})$ at the cell edges
$\dot{l}_{\pm}=\frac{1}{\xi}\partial_{x}\sigma(l_{\pm}(t),t)+v_{p}.$ (16)
This corresponds to symmetric polymerization where actin filaments are extended at the leading and retracted at the trailing edge with the same speed. In order to keep the total mass of myosin constant, polymerization requires novel flux boundary conditions
$\partial_{x}c_{A,B}(l_{\pm},t)=-\frac{v_{p}}{D}\frac{\left(c_{\infty}-c_{A,B}(l_{\pm},t)\right)^{2}}{c_{A,B}(l_{\pm},t)}.$ (17)
This change poses difficulties when formulating the equations for steady-state motility as additional boundary terms are required. In the following analysis, we focus on the steady-state equations previously derived under no-flux boundary conditions. As a result, there is a slight loss of total myosin mass; however, this effect remains negligible for small polymerization velocities and over short time scales. While this approach is approximate, it nonetheless offers valuable qualitative insight into the system’s behavior.
Fig. 7(a) and (b) show the bifurcation analysis in steady state subject to small polymerization velocities. Any nonzero polymerization turns the pitchfork bifurcation into an imperfect bifurcation, which means that it splits into a biased motile solution branch, a stable solution branch with opposite polymerization- and contraction-driven polarizations, and an unstable branch unpolarized in the myosin distribution except for retrograde flow from polymerization. This change in bifurcation type however neither changes the polarization mechanism from contraction nor the qualitative isoform distribution (cf. Fig. 7 (c)).
@@ IV Discussion
The two-species model introduced here recovers the experimental distributions of the isoforms and in addition predicts the existence of a variety of migratory modes. Regarding the steady states, in the non-motile steady state both isoforms are homogeneously distributed and cause a symmetric active contractile stress. At sufficiently large Péclet numbers, however, this symmetry is broken and polarized concentration fields and, consequently, polarized active stress profiles emerge. This onset of polarization and migration is due to a supercritical pitchfork bifurcation. For a steadily moving cell beyond the bifurcation, we find the experimentally observed, isoform-specific myosin distribution profiles [12]: the faster NMIIA accumulates at the leading edge, while the stronger NMIIB localizes at the trailing edge. We also showed that our model can be extended to describe the effect of uneven actin polymerization at the boundaries, which immediately stabilizes motile solutions and leads to an imperfect pitchfork bifurcation, but otherwise does not change our conclusions.
Our one-dimensional model is motivated by the recent body of experimental work of cells migrating along lines and in channels. In order to analyze higher dimensions, other model classes such as phase field models are easier to implement [8, 56], although a recent mathematical analysis showed how a rigorous bifurcation analysis can be performed on a 2D model similar to the one described here [43]. Due to the similarities of this work with the one-dimensional one-species model [21], we would also expect our two-species model to behave similarly in two dimensions.
@@
We have explored further modes of motility besides the steady state solutions. Oscillations occur in case that the more contractile isoform is also more diffusive (i.e. $P_{B}>P_{A},\ K_{r}>1$), via a cyclic pull-push-mechanism. This mechanism works best for similar total abundances of both isoform, $c_{r}\sim 1$. It has been suggested that different myosin isoforms can be regulated independently [57, 58, 59], which would allow for the relative diffusion to be controlled via $K_{r}$. Albeit not observed in the animal cell type investigated in Refs. [16, 12], it is not unlikely that these oscillations might occur in other cell types (e.g. from other species). Another interesting option is to genetically engineer cells such that these oscillations occur, either spontaneously or triggered by external control, in particular by optogenetics. If oscillations can be controlled in this manner, it also would be easy to distinguish them from oscillations that occur as a result of non-linear friction, as described by clutch models [60, 61, 62].
@@
In the oscillatory state, the amplitude of the cell length oscillation is much smaller than the one of velocity. This suggested to study the rigid limit, which enabled an analytic linear stability analysis and allowed to determine the full state diagram of the system, containing sessile cells, cells with transient oscillations and persistent oscillations, as well as steady motile cell solutions. This phase diagram is in good agreement with finite volume simulations of the full model.
@@
In summary, here we have shown that active gel theory can be extended to also describe the effect of having two complementary isoforms at work. The fact that we could derive the same model from both kinetic and a thermodynamic perspectives demonstrate its fundamental nature and paves the way to also include other biologically relevant effects into the model. In particular, a third major molecular constituent of contractile structures in cells is $\alpha$-actinin, which like NMII also exists in different isoforms and which also competes for binding to the actin filaments. In our model framework, such a competitor would increase excluded volume for the NMII and therefore decrease active stress. At the same time, $\alpha$-actinin might increase elasticity and therefore make the cell effectively stiffer. It would be interesting to include such effects, but this also would introduce many more model parameters, that then had to be estimated from experiment. The same is true for replacing the simple friction by a molecular clutch [60, 63]. Despite this complexity, however, the work presented here clearly shows that molecular details can be fruitfully included in fundamental models for cell motility.
@@ V Acknowledgments
Appendix A Parameter Analysis
In the following, we analyze the system’s dependence on several parameters and their effect on the dynamics. We use steady state solutions throughout. For all analyzed values we find supercritical pitchfork bifurcations at $\mathcal{K}_{crit}$ in analogy to what is shown in Fig. 2(a).
Figure S1: Concentration profiles for different saturation concentrations $c_{\infty}$ and relative diffusivities $K_{r}$. All profiles taken at $\mathcal{K}=1000$, contractilities $P_{A}=P_{B}=0.06$ and total concentration ratio $c_{r}=1$. For $K_{r}=1$ both profiles are equal.
Fig. S1 shows the effect of various combinations of parameter values of saturation concentration $c_{\infty}$ and relative diffusivity $K_{r}$. Following the rows, we see that a low saturation concentration flattens and widens the concentration profiles. In contrast, high saturation allows for sharper peaks and overall spatially narrower profiles inside the cell. This holds for all $K_{r}$. Larger saturation concentrations allow for stronger polarization at the edges. Hence, the bifurcation point $\mathcal{K}_{crit}$ marking the transition from sessile to motile states shifts to smaller values for increasing saturation concentration. The columns reveal the influence of relative diffusion. The isoform with the lowest diffusivity sits further in the back (species $B$ for $K_{r}<0$, species $A$ for $K_{r}>0$). We can explain this by the transport of myosin along the retrograde actin flow. The retrograde flow moves towards the trailing edge. For the cases depicted here of positive cell velocity we find a flow towards the left. During advection, the motors also diffuse away and unbind from actin. The higher the diffusion, the less motors arrive at the trailing edge to initiate motility. This also explains why the critical Péclet number ($\mathcal{K}_{crit}$) increases with increasing $K_{r}$. According to Eq. (10), the diffusion of species $A$ is not affected by $K_{r}$, which only influences the diffusion of species $B$. Hence, an increase in $K_{r}$ leads to a higher total diffusion with less biased concentration and therefore less asymmetric active stress profiles, causing a later onset of motility. Once the space at the trailing edge is occupied by the less diffusive species, the other species can only place itself closer to the cell center due to the volume exclusion interaction.
Figure S2: Concentration profiles and active stress for varying relative contractility and concentration ratios. All profiles are plotted at $\mathcal{K}=2500$. Fixed parameters are $K_{r}=1.0,c_{\infty}=5$. The following values for the respective contractility were used: first row: $P_{A}=0.02,P_{B}=0.01$; second row: $P_{A}=P_{B}=0.015$; third row: $P_{A}=0.01,P_{B}=0.02$.
[…]